# Grammar Induction from Visual, Speech and Text

### v1

@@ Abstract
Grammar Induction could benefit from rich heterogeneous signals, such as text, vision, and acoustics. In the process, features from distinct modalities essentially serve complementary roles to each other. With such intuition, this work introduces a novel unsupervised visual-audio-text grammar induction task (named VAT-GI), to induce the constituent grammar trees from parallel images, text, and speech inputs. Inspired by the fact that language grammar natively exists beyond the texts, we argue that the text has not to be the predominant modality in grammar induction. Thus we further introduce a textless setting of VAT-GI, wherein the task solely relies on visual and auditory inputs. To approach the task, we propose a visual-audio-text inside-outside recursive autoencoder (VaTiora) framework, which leverages rich modal-specific and complementary features for effective grammar parsing. Besides, a more challenging benchmark data is constructed to assess the generalization ability of VAT-GI system. Experiments on two benchmark datasets demonstrate that our proposed VaTiora system is more effective in incorporating the various multimodal signals, and also presents new state-of-the-art performance of VAT-GI.
@@
[…]
@@
1 Introduction
Human language knowledge involves various aspects, such as the vocabulary, phonetics, morphology, syntax, semantics and pragmatics of languages [5, 37, 2, 9]. Among these, inferring the underlying grammatical rules and linguistic patterns of a language plays a significant role in language learning. The process is also tasked as grammar induction (GI) [25, 56, 45, 13] , aiming to uncover the latent structure of language constituents from natural inputs in an unsupervised manner.
In the community, the mainstream explorations on GI pay an extensive focus on the textual signals [8, 24, 49, 4, 18, 23, 59, 46], as texts are the major medium of language. However, it is important to note that human always acquire language knowledge with multimodal signals, i.e., text, vision and acoustic. Knowing this fact, researchers carry out GI using different modalities of information, such as vision-language grammar learning [59, 15, 54, 47, 48], speech-based textless GI [32] and video-aided GI [58, 57, 39, 42].
Unfortunately, most existing researches on GI tend to overlook the mutually complementary contributions of the information in broad-covering modalities, which together can play crucial roles in the acquisition of language throughout phylogenetic development. This can be exemplified by the cases in Fig. 1: the pixel regions in a visual image are often associated with the noun phrase structure, such as ‘a woman’ and ‘a tour map’, while speech intonation and rhythm help segment sentences from a constituent-level perspective, such as ‘scanning a tour map’ and ‘with her smartphone’. This thus motivates us to introduce a novel task of unsupervised visual-audio-text grammar induction (VAT-GI), with the aim of exploring the language learning of NLP related AI systems in realistic scenario. Particularly, VAT-GI operates over parallel images, audios and texts, and extracts the shared latent constituency structures that encompass all these modalities (see §$\S$2 for task definition). As depicted in Fig. 1, the terminal nodes of the tree are the individual words, while non-terminal nodes are notional constituent spans, where nodes also associate with visual regions and speech clips. Different sources of modalities together form an integral constituency grammar structure.
@@ 1 Introduction
  • Text: Language intrinsically encompasses structural information by virtue of its inherent compositionality [5, 35, 44], thereby providing a straightforward basis for grammar induction.
@@
  • Vision: The smaller spatial regions of pixels combine to form larger regions with higher-level visual semantics [15, 59, 54], which correspond to the hierarchical structure observed among textual constituent spans.
@@
  • Audio: Speech conveys structural information through various intonation and rhythm patterns, naturally depicting the discontinuity and continuity of phonemes [41, 50, 3, 36, 14].
@@
With the aforementioned observations, in this work we present a nichetargeting framework to approach VAT-GI. Built upon the deep inside-outside recursive autoencoder (Diora) model [8] for effective text-based GI, we devise a novel visual-audio-text Diora (namely VaTiora), which extends the Diora with the further capability of capturing specific structural characteristics of three modalities of sources, such as the compositionality of texts, the hierarchical spatial regions of images and the rhythm patterns of speech, and integrating them properly. As shown in Fig. 3, first, rich modal-specific and complementary features, such as text embedding, visual regions and voice pitch, etc., are constructed from the input text, image and audio sources, respectively, all of which are effectively integrated into the inside-outside recursive autoencoding. Then, in the feature interaction phase, text embeddings are first fused with audio features via cross-attention to obtain the token-wise text-audio representations for constructing the span vectors; these span vectors are further mapped with fine-grained visual region features by an attention-aware fusion, thereby enhancing the span representations. Moreover, the composition probabilities of visual regions are considered when calculating the span scores, during which the intonation pitch frequency features and voice activity features (representing the rhythm pattern) are also incorporated. Through such cohesive integration, the cross-modal signals mutually reinforce each other for more accurate constituency parsing in the inside-outside pass of VaTiora. We follow [8] to train the VaTiora with structure reconstruction and contrastive learning. In addition, we further introduce a cosine similarity learning objective across three types of features for representation alignment, bridging the heterogeneity among modalities. Additionally, we contend that text is not a fundamental modality for VAT-GI, as supported by the physical fact that language grammar natively exists beyond the texts. Hence, we introduce a textless setting for VAT-GI, wherein the task solely relies on visual and auditory inputs, and the constituent tree is no longer structured around individual words but instead the segmented speech clips. Accordingly, we specially propose the aligned span-clip F1 metric (SCF1) to measure the textless GI, due to the inaccessibility of golden word segments of speeches in this setting.
@@
VAT-GI can be evaluated on the public SpokenCOCO dataset [17] with well-aligned images, texts and audio speeches. However, SpokenCOCO has inherent limitations in terms of monotonous styles of visual scenes and speech tones, as well as short captions, which largely weakens the model generalization ability to complex signal inputs. To combat that, we present a more challenging test set for VAT-GI, namely SpokenStory, which advances in richer visual scenes, timbre-rich audio and deeper constituent structures. We collect 1,000 aligned image-text from Localized Narratives [43], and further record speech for each sentence by human speakers. The experimental results on the datasets suggest that leveraging multiple modality features helps better grammar induction. Further analyses demonstrate that VaTiora is effective in incorporating the various multimodal signals, and yielding new state-of-the-art induction performance of VAT-GI task.
@@
  • We for the first time introduce an important yet challenging task, unsupervised visual-audio-text grammar induction, to emulate the phylogenetic language acquisition;
@@
  • We present a novel VAT-GI framework, VaTiora, to properly navigate multimodal input sources;
  • We also propose a new metric for the textless VAT-GI evaluation;
@@
  • We newly contribute a challenging benchmark data, SpokenStory, for VAT-GI;
  • New state-of-the-art performance on the benchmarks are shown. Our resources will be open to facilitate the community.
@@ 2 Task Definition
Formulation. Given a sentence $X=\{x_{1},x_{2},...,x_{n}\}$ with $n$ words, an associated image $I$ and a piece of speech $S$ for $X$, the goal of unsupervised VAT-GI is to induce a constituency tree structure $T$ from $X$, $I$ and $S$ without any supervision of constituent structure annotations for training. As shown in Fig. 1, the output tree structure $T$ is formed in a phrase constituent tree by Chomsky Normal Form (CNF) [6], where each non-terminal node in the tree has exactly two children. Each node in $T$ contains a text span $(i,j),1\leq i\leq j\leq n$, a region box $b_{i,j}\in\mathbb{R}^{4}$ in $I$ and a speech clip $c_{i,j}\in\mathbb{R}^{2}$, where the text span $(i,j)$ is grounded to the region box $b_{i,j}$ and aligned to the speech clip $c_{i,j}$. For the textless setting, the inputs of VAT-GI are only $I$ and $S$. Thus the node set of output $T$ contains only $c_{i,j}$ and $b_{i,j}$.
Evaluation. We adopt two widely-used metrics, averaged corpus-level F1 (Corpus-F1) and averaged sentence-level F1 (Sent-F1), following [24]. Corpus-level F1 calculates precision/recall at the corpus level to obtain F1, while sentence-level F1 calculates F1 for each sentence and averages across the corpus. In this paper, we follow [24] discard trivial spans (length < 2) and evaluate on sentence-level F1 per recent work. In textless settings, where there is no text script for speech-to-text alignment, we introduce a new metric, aligned span-clip F1 (SCF1), to measure the quality of the textless constituency tree. SCF1 first aligns predicted speech clips to ground-truth speech clips via temporal IoU [38]. We use a greedy mapping from start to end, aligning predicted clip and ground-truth segmented clip if their tIoU is over a predefined threshold $p$. The tIoU [38] is a metric to measure the overlap ratio on the temporal dimension. If a ground-truth clip has multiple mapped predictions, we keep the one with the largest tIoU. After alignment, we then calculates the F1 score based on TP, FP and FN:
$\begin{split}\text{TP}&=\text{Count}(\text{tIoU}(pred,gt)\geq p),\\ \text{FP}&=\text{Count}(\text{tIoU}(pred,gt)<p\text{ or }pred\notin gt),\\ \text{FN}&=\text{Count}(\text{tIoU}(pred,gt)<p\text{ or }gt\notin pred),\end{split}$ (1)

### v2

# Programming, Writing & Editing

@@ Abstract
Grammar Induction (GI) seeks to uncover the underlying grammatical rules and linguistic patterns of a language, positioning it as a pivotal research topic within Artificial Intelligence (AI). Although extensive research in GI has predominantly focused on text or other singular modalities, we reveal that GI could significantly benefit from rich heterogeneous signals, such as text, vision, and acoustics. In the process, features from distinct modalities essentially serve complementary roles to each other. With such intuition, this work introduces a novel unsupervised visual-audio-text grammar induction task (named VAT-GI), to induce the constituent grammar trees from parallel images, text, and speech inputs. Inspired by the fact that language grammar natively exists beyond the texts, we argue that the text has not to be the predominant modality in grammar induction. Thus we further introduce a textless setting of VAT-GI, wherein the task solely relies on visual and auditory inputs. To approach the task, we propose a visual-audio-text inside-outside recursive autoencoder (VaTiora) framework, which leverages rich modal-specific and complementary features for effective grammar parsing. Besides, a more challenging benchmark data is constructed to assess the generalization ability of VAT-GI system. Experiments on two benchmark datasets demonstrate that our proposed VaTiora system is more effective in incorporating the various multimodal signals, and also presents new state-of-the-art performance of VAT-GI. Further in-depth analyses are shown to gain a deep understanding of the VAT-GI task and how our VaTiora system advances. Our code and data: https://github.com/LLLogen/VAT-GI/
@@
1 Introduction
Within the field of AI, human language acquisition is one of the important research topics. Human language knowledge involves various aspects, such as the vocabulary, phonetics, morphology, syntax, semantics and pragmatics of languages Chomsky (1957); Osherson and Gleitman (1995); Briscoe (2000); Flynn and O’Neil (2012); Fei et al. (2020, 2022, 2021, 2024); Chen et al. (2024); Jia et al. (2024). Among these, inferring the underlying grammatical rules and linguistic patterns of a language plays a significant role in language learning. The process is also tasked as grammar induction Kim et al. (2019b); Yogatama et al. (2017); Shen et al. (2019); Harwath et al. (2018), aiming to uncover the latent structure of language constituents from natural inputs in an unsupervised manner. In the community, the mainstream explorations on GI pay an extensive focus on the textual signals Drozdov et al. (2019); Kim et al. (2019a); Spitkovsky et al. (2010); Choi et al. (2018); Hu et al. (2022); Kim et al. (2020); Zhao and Titov (2020); Shi et al. (2020), as texts are the major medium of language. However, it is important to note that human always acquire language knowledge with multimodal signals, i.e., text, vision and acoustic. Knowing this fact, researchers carry out GI using different modalities of information, such as vision-language grammar learning Zhao and Titov (2020); Hong et al. (2021); Wan et al. (2022); Shi et al. (2019, 2018), speech-based textless GI Lai et al. (2023a) and video-aided GI Zhang et al. (2021, 2022); Perdani (2022); Piergiovanni et al. (2020).
Unfortunately, most existing researches on GI tend to overlook the mutually complementary contributions of the information in broad-covering modalities. Actually, information and signals from all different modalities together can play crucial complementary roles in the acquisition of language throughout phylogenetic development. This can be exemplified by the cases in Figure 1: the pixel regions in a visual image are often associated with the noun phrase structure, such as ‘a woman’ and ‘a tour map’, while speech intonation and rhythm help segment sentences from a constituent-level perspective, such as ‘scanning a tour map’ and ‘with her smartphone’. This thus motivates us to introduce a novel task of unsupervised visual-audio-text grammar induction (VAT-GI), with the aim of exploring the language learning of NLP-related AI systems in realistic scenarios. Particularly, VAT-GI operates over parallel images, audios and texts, and extracts the shared latent constituency structures that encompass all these modalities (see §$\S$3 for task definition). As depicted in Figure 1, the terminal nodes of the tree are the individual words, while non-terminal nodes are notional constituent spans, where nodes also associate with visual regions and speech clips. Different sources of modalities together form an integral constituency grammar structure.
@@ 1 Introduction
  • Text: Language intrinsically encompasses structural information by virtue of its inherent compositionality Chomsky (1957); Mikolov et al. (2013); Shen et al. (2018), thereby providing a straightforward basis for grammar induction.
@@
  • Vision: The smaller spatial regions of pixels combine to form larger regions with higher-level visual semantics Hong et al. (2021); Zhao and Titov (2020); Wan et al. (2022), which correspond to the hierarchical structure observed among textual constituent spans.
@@
  • Audio: Speech conveys structural information through various intonation and rhythm patterns, naturally depicting the discontinuity and continuity of phonemes Pickett (1999); Stockwell (1960); Byun and Tessier (2016); Müller (2002); Heinz (2011).
@@
With the aforementioned observations, in this work we present a nichetargeting framework to approach VAT-GI. Built upon the deep inside-outside recursive autoencoder (Diora) model Drozdov et al. (2019) for effective text-based GI, we devise a novel visual-audio-text Diora (namely VaTiora), which extends the Diora with the further capability of capturing specific structural characteristics of three modalities of sources, such as the compositionality of texts, the hierarchical spatial regions of images and the rhythm patterns of speech, and integrating them properly. As shown in Figure 3, first, rich modal-specific and complementary features, such as text embedding, visual regions and voice pitch, etc., are constructed from the input text, image and audio sources, respectively, all of which are effectively integrated into the inside-outside recursive autoencoding. Then, in the feature interaction phase, text embeddings are first fused with audio features via cross-attention to obtain the token-wise text-audio representations for constructing the span vectors; these span vectors are further mapped with fine-grained visual region features by an attention-aware fusion, thereby enhancing the span representations. Moreover, the composition probabilities of visual regions are considered when calculating the span scores, during which the intonation pitch frequency features and voice activity features (representing the rhythm pattern) are also incorporated. Through such cohesive integration, the cross-modal signals mutually reinforce each other for more accurate constituency parsing in the inside-outside pass of VaTiora. We follow Drozdov et al. (2019) to train the VaTiora with structure reconstruction and contrastive learning. In addition, we further introduce a cosine similarity learning objective across three types of features for representation alignment, bridging the heterogeneity among modalities.
Additionally, we introduce a textless setting for VAT-GI, wherein the task solely relies on visual and auditory inputs, and the constituent tree is no longer structured around individual words but instead the segmented speech clips. The motivation for setting the textless VAT-GI comes from the consideration of the languages that lack written forms, e.g., the languages of minority ethnic groups in certain regions. We contend that text is not a fundamental modality for VAT-GI, as supported by the physical fact that language grammar natively exists beyond the texts. Through investigating the feasibility of textless VAT-GI, we uncover the potential of modeling syntactic structures using non-textual modalities. Accordingly, we specially propose the aligned span-clip F1 metric (SCF1) to measure the textless VAT-GI, due to the inaccessibility of golden word segments of speeches in this setting.
@@
VAT-GI can be evaluated on the public SpokenCOCO dataset Hsu et al. (2021) with well-aligned images, texts and audio speeches. However, SpokenCOCO has inherent limitations in terms of monotonous styles of visual scenes and speech tones, as well as short captions, which largely weakens the model generalization ability to complex signal inputs. To combat that, we present a more challenging test set for VAT-GI, namely SpokenStory, which advances in richer visual scenes, timbre-rich audio and deeper constituent structures. We collect 1,000 aligned image-text from Localized Narratives Pont-Tuset et al. (2020), and further record speech for each sentence by human speakers. The experimental results on the datasets suggest that leveraging multiple modality features helps better grammar induction. Further analyses demonstrate that VaTiora is effective in incorporating the various multimodal signals, and yielding new state-of-the-art induction performance of VAT-GI task.
@@
  • We for the first time introduce an important yet challenging task, unsupervised visual-audio-text grammar induction, to better emulate the human-level phylogenetic language acquisition;
@@
  • We present a novel VAT-GI model, VaTiora, to properly navigate multimodal input sources;
@@
  • We newly contribute a challenging benchmark data, SpokenStory, for VAT-GI; and also propose a new metric for the textless VAT-GI evaluation;
  • Our proposed method sets a strong-performing benchmark result for follow-up research. Our resources will be open to facilitate the community.
The rest of the paper is organized as follows. The next section contains the background and related work of textual grammar induction and multi-modal grammar induction. After that, in section 3 we present the definition visual-audio-text grammar induction task and how to evaluate it. In section 4, we elaborate on the details of the proposed VaTiora framework, which incorporates multiple cues from three modalities into an inside-outside algorithm. Following that, in the experiment section, we compare our methods with strong grammar induction baselines, and explore how our method advances via in-depth quantitative and qualitative analysis. Finally, we shed the light on the future work of this research, and then conclude the paper.
2 Related Work
In the field of AI, research draws inspiration from language to enhance the development of more intelligent systems and models. Among all topics, the task of grammar induction aims to infer the underlying grammatical rules and linguistic patterns of a language, thereby gaining a deeper understanding of the principles governing human intelligence. GI unsupervisedly determines the constituent syntax of language, in the format of a phrase-structure tree, which intuitively depicts the language compositionality, i.e., how the small components (e.g., words or phrases) combine to larger constituents (e.g., clauses or sentences). Such a process essentially emulates the human language acquisition. Thus, GI Lari and Young (1990) has long been a fundamental topic in natural language processing (NLP) Cohen and Smith (2008); Spitkovsky et al. (2010); Klein and Manning (2002); Hu et al. (2022); Kim et al. (2020); Zhao and Titov (2020); Hong et al. (2021); Wan et al. (2022); Shi et al. (2019).
According to the categories of information modalities, all existing GI research can be categorized into text-based GI, vision-language-based GI, and speech-based GI. Below, we survey relevant studies within each of these groups.
Textual Grammar Induction. The majority of previous GI studies pay the focus on the language domain with textual corpora Cohen and Smith (2008); Lari and Young (1990); Spitkovsky et al. (2010); Klein and Manning (2002); Hu et al. (2022); Kim et al. (2020). Generally, there are two types of frameworks for text-only grammar induction. The first is PCFG-based methods, which assume the context-free grammar rules exist with a probability and estimates the likelihood of each latent rule Kim et al. (2019a); Kurihara and Sato (2006); Wang and Blunsom (2013). One way of inducing this probabilistic grammar is to fix the structure of a grammar and then find the set of optimal parameters such that the resulting grammar best explains the language, which is usually approximated by a training corpus. Early works induce PCFG via statistic-based methods, such as EM algorithm Lari and Young (1990). Following efforts focus on improving it by carefully-crafted auxiliary objectives Klein and Manning (2002), priors or non-parametric models Wang and Blunsom (2013), and manually-engineered features Golland et al. (2012). Kim et al. (2019a) first propose to parameterize the PCFG’s rule probability with a trainable neural network. Jin et al. Jin et al. (2018) applies depth bounds within a chart-based Bayesian PCFG inducer, limiting the search space of an unsupervised PCFG inducer. The other type of framework is autoencoder-based methods, which estimate span composition likelihood without assuming context-free grammar. Drozdov et al. (2019) first propose the Diora model, which incorporates the inside-outside algorithm into a latent tree chart parser.
Visual-Language Grammar Induction. The visually-grounded GI has received increasing research attention Zhao and Titov (2020); Hong et al. (2021); Wan et al. (2022); Shi et al. (2019). Since visual regions (e.g., objects of interest) intuitively encompass the correspondence of the textual spans (e.g., noun phrases), the visions are imported as a type of enhanced signal for better constituency tree parsing. Shi et al. (2019) first propose to import visual features into grammar induction, mapping visual and text via visual-semantic embedding. The following works extend text-only GI models with extra visual information, such as VC-PCFG Zhao and Titov (2020), VG-NSL Shi et al. (2019) and Cliora Wan et al. (2022). The visually grounded compound PCFGs (VC-PCFG) extends the compound PCFG model (C-PCFG) by including a matching model between images and text. Similar performance gain is also observed with VG-NSL. Instead of fusing information of the entire image into the phrase representations, Cliora uses region-based fine-grained alignment to have a thorough understanding of the image. There are also video-aided GI, where, compared with a single image, video frames can further facilitate the understanding of motion dynamics for the correspondence of textual predicates Zhang et al. (2021, 2022); Perdani (2022); Piergiovanni et al. (2020). Zhang et al. Zhang et al. (2021, 2022) first propose video-aided GI models that prove videos can provide even richer information than images for grammar induction, including not only static objects but also actions and state changes useful for inducing verb phrases.
Speech-oriented Grammar Induction. On the other hand, acquiring languages from speech has also gained consistent research interests Klasinas et al. (2013); Iosif et al. (2018); Jansen et al. (2013); Harwath et al. (2020, 2018), where the acoustic-prosodic features (e.g., phonetics, phonology) can offer important clues for the syntax induction from different perspectives than the visions and texts Byun and Tessier (2016); Müller (2002); Heinz (2011). Lai et al. Lai et al. (2023b) first attempt incorporate audio signals for grammar induction and present the novel AV-NSL learner. They segment the speech waveform into sequences of word segments, and subsequently induce phrase structure using the inferred segment-level continuous representations. In this work, we take a combined holistic viewpoint, and investigate the GI under multimodal information sources, i.e., introducing a novel visual-audio-text grammar induction task. This can be quite intuitive, as we humans always perceive the world with varied sensory inputs that can partially share the same common structures and meanwhile preserve distinctly complementary features, which together help achieve more effective GI.
Despite the significant achievements and extensive attention garnered by prior GI research, current approaches remain confined to single modalities, such as text or others, often overlooking the mutually complementary contributions of information from diverse modalities. In reality, information and signals from various sources—including text, vision, and acoustics—collectively and complementarily play crucial roles in language acquisition throughout phylogenetic development. Leveraging rich heterogeneous signals could greatly benefit GI. To address this, we propose a novel task, unsupervised visual-audio-text grammar induction (VAT-GI), to emulate real human-level phylogenetic language acquisition.
@@ 3 Task Definition
Formulation. Given a sentence $X=\{x_{1},x_{2},...,x_{n}\}$ with $n$ words, an associated image $I$ and a piece of speech $S$ for $X$, the goal of unsupervised VAT-GI is to induce a constituency tree structure $T$ from $X$, $I$ and $S$ without any supervision of constituent structure annotations for training. As shown in Figure 1, the output tree structure $T$ is formed in a phrase constituent tree by Chomsky Normal Form (CNF) Chomsky (1959), where each non-terminal node in the tree has exactly two children. Each node in $T$ contains a text span $(i,j),1\leq i\leq j\leq n$, a region box $b_{i,j}\in\mathbb{R}^{4}$ in $I$ and a speech span $c_{i,j}\in\mathbb{R}^{2}$, where the text span $(i,j)$ is grounded to the region box $b_{i,j}$ and aligned to the speech span $c_{i,j}$. For the textless setting, the inputs of VAT-GI are only $I$ and $S$. Thus the node set of output $T$ contains only $c_{i,j}$ and $b_{i,j}$.
@@
Evaluation. We adopt two widely-used metrics, averaged corpus-level F1 (Corpus-F1) and averaged sentence-level F1 (Sent-F1), following Kim et al. (2019a). Corpus-level F1 calculates precision/recall at the corpus level to obtain F1, while sentence-level F1 calculates F1 for each sentence and averages across the corpus. In this paper, we follow Kim et al. (2019a) discard trivial spans (length < 2) and evaluate on sentence-level F1 per recent work. In textless settings, where there is no text script for speech-to-text alignment, we introduce a new metric, aligned span-clip F1 (SCF1), to measure the quality of the textless constituency tree. Suppose the output of the textless VAT-GI contains a sequence of speech clips $\{clip^{p}_{m}\}$. The predicted constituency tree is denoted as the pairs of speech clips $\{c^{p}\}=\{(clip^{p}_{h},clip^{p}_{t})\}$, where each clip represents a token. SCF1 first aligns predicted speech clips to ground-truth speech clips via tIoU (Temporal Intersection over Union) Peng and Harwath (2022), which represents the overlap between two time intervals. It is a metric to measure the overlap ratio on the temporal dimension. We use a greedy mapping from start to end, aligning predicted clip and ground-truth segmented clip if their tIoU is over a predefined threshold $p$. In details, we traverse all the ground-truth clips, denoting $\{clip^{g}_{n}\}$, to search the aligned clips. For each $clip^{g}_{n}$, we calculate the tIoU between $clip^{g}_{n}$ with all the predicted clips $\{clip^{p}_{m}\}$. A predicted clip $clip^{g}_{n}$ is considered as the aligned one with $clip^{p}_{m}$ only when tIoU$(clip^{g}_{n},clip^{p}_{m})>p$. If a ground-truth clip has multiple mapped predictions, we keep the one with the largest tIoU. After alignment, we then calculate the F1 score based on TP, FP and FN. We say the two spans are “matched” when both their head and tail clips are aligned respectively. Let $\{c^{g}\}=\{(clip^{g}_{h},clip^{g}_{t})\}$ be the ground-truth spans, then we have:
$\begin{split}\text{TP}&=\text{Count}(pred\text{ and }gt\text{ are matched},pred\in\{c^{p}\}\text{ and }gt\in\{c^{g}\}),\\ \text{FP}&=\text{Count}(pred\in\{c^{p}\}\text{ that no }gt\in\{c^{g}\}\text{ could match}),\\ \text{FN}&=\text{Count}(gt\in\{c^{g}\}\text{ that no }pred\in\{c^{p}\}\text{ could match}),\end{split}$ (1)